OpenStudy (ac3):

use equation 1 to find a power series representation for f(x)=ln(1-x). What is the radius of convergence?

OpenStudy (ac3):

\[f(x)=\ln(1-x)\]

OpenStudy (ac3):

@Vocaloid Any idea?

OpenStudy (anonymous):

Not sure what you mean by equation (1), but I'll take a wild guess and suppose it's \[\frac{1}{1-x}=\sum_{n=0}^\infty x^n\quad\text{for }|x|<1\] Notice that \[\frac{d}{dx}\ln(1-x)=-\frac{1}{1-x}=-\sum_{n=0}^\infty x^n\] Integrating, you have \[\ln(1-x)=-\sum_{n=0}^\infty \frac{x^{n+1}}{n+1}+C\] If we consider \(x=0\), you would find that \[\ln(1-0)=-\sum_{n=0}^\infty \frac{0^n}{n+1}+C~\implies~C=0\]

OpenStudy (ac3):

Alright so we have \[\sum_{0}^{\infty} \frac{ x ^{n+1} }{ n+1 }\]

OpenStudy (ac3):

using that do we now we use that to find the power series of xln(1-x)

OpenStudy (ac3):

@SithsAndGiggles Would I just multiply the whole thing by x getting. \[\sum_{n=0}^{\infty} \frac{ x ^{n+2} }{ n+1 }\]

OpenStudy (ac3):

??

OpenStudy (anonymous):

Right, if the power series for \(f(x)\) is given by \(S\), then the power series for \(x f(x)\) is \(xS\). You're missing the minus sign, btw.

OpenStudy (ac3):

THANK YOU!!!

OpenStudy (anonymous):

yw

OpenStudy (ac3):

hold on though.

OpenStudy (ac3):

we're on the last part of the entire question so now by putting x=1/2 in your result from part a (that's our first one), express ln2 as te sum of an infintie series.

OpenStudy (anonymous):

Okay, so if \(x=\dfrac{1}{2}\), then \[\ln\left(1-\frac{1}{2}\right)=\ln\frac{1}{2}=\ln2^{-1}=-\ln2\]

OpenStudy (ac3):

That's it?

OpenStudy (anonymous):

Yep! The series representation is simple enough, you're just replacing \(x\) with \(\dfrac{1}{2}\). \[-\ln2=-\sum_{n=0}^\infty \frac{\left(\frac{1}{2}\right)^{n+1}}{n+1}\]which you can rewrite in several ways.

OpenStudy (ac3):

Thank you dude your freaking awesome!!!

OpenStudy (anonymous):

yw, here to help :)